Suppose the base and height of a triangle sum to 12, and the area of the triangle is 16 square cm. What are the dimensions (base and height) of the triangle?
Please break the problem down and show me how you solved it,
Thank you!

Answer by mananth(16946)

let base =b and height =h
(b+h)=12
area of triangle = (1/2) * b*h
(1/2) bh = 16
bh =16*2=32
h= 32/b
Now (b+h)=12
(b+32/b ) = 12
multiply equation by b
b^2+32=12b
b^2-12b+32=0
b^2-8b-4b+32=0
b(b-8)-4(b-8)=0
(b-8)(b-4)=0
b=4 OR 8
if b=4 h=8
if b=8 h= 4